OpenStudy (anonymous):

Solve the equation: x^2+18x+81=25

OpenStudy (uri):

x^2+18x+81=25 x^2+18x=25-81 x^2+18x=-56 x^2+18x+56 =0

OpenStudy (uri):

Now do we have to use Quad formula?

OpenStudy (anonymous):

18^2-4*56=324-224=100 ,then delta =10 , x1=-18+10/2=-4 , x2 =-18-10/2=-14